OpenStudy (lilai3):

@hartnn please help: 3) 7.4 km squared equals to how many meters squared 5) 3 yd squared equals to how many inches squared 7) 0.03( with a bar over the three only) miles squared equals to how many yd squared?

hartnn (hartnn):

how many metres are there in a km ?

OpenStudy (campbell_st):

1000 m = 1km so 1km^2 = 1000 x 1000 or 1 000 000 m^2 so if you have 7.4 km^2 = 7.4 x 1 000 000

OpenStudy (lilai3):

this is really random but @hartnn are you from britain or europe by any chance

OpenStudy (lilai3):

cuz u spelled meters wrong

hartnn (hartnn):

hmm...

OpenStudy (lilai3):

lols anyway back to the subject...

hartnn (hartnn):

no.

OpenStudy (lilai3):

1000******

hartnn (hartnn):

yes.

hartnn (hartnn):

so, km squares = how many meter squared ?

OpenStudy (lilai3):

uhm...100,000? hm...to be honest with you, i don't really understand the squaring part please explain you are a good person:) i'm sorry i don't use periods :P

hartnn (hartnn):

1 000^2 = 1 000 000 so 7.4 km squared = ?? meter squared ??

OpenStudy (lilai3):

am i rite tho? i don't get the ^ signs... ???

hartnn (hartnn):

no, \(1000^2=1000000\)

hartnn (hartnn):

if there are 3 xeros before squaring, there will be 6 zeros after squaring.

hartnn (hartnn):

*zeros

OpenStudy (lilai3):

oh i get it you add thtrree more zeros not squaee the rhree zeros rigte

hartnn (hartnn):

hahaha..... no problem

hartnn (hartnn):

did u get that ?

OpenStudy (lilai3):

sorat yeah

hartnn (hartnn):

1 000^2 = 1 000 000 so 7.4 km squared = ?? meter squared ??

OpenStudy (lilai3):

740000000?

hartnn (hartnn):

7.4 * 1000000 = 7400000

OpenStudy (lilai3):

oh.oops. so that the answer?

hartnn (hartnn):

yup.

OpenStudy (lilai3):

i have another question : 2.2 yd cubed equals to how many ft cubed?